1.      An individual seller’s monthly supply of downloadable e-books is given by the equation

Qseb = −6.45 3.7Peb – 7.5W

where Qseb is number of e-books supplied, Peb is the price of e-books in euros, and W is the wage rate in euros paid by e-book sellers to laborers. Assume that the price of e-books is €10.68 and wage is €10. The supply side of the market consists of a total of eight identical sellers in this competitive market.

a.      Determine the market aggregate supply function.

b.     Determine the inverse market supply function.

Determine the slope of the aggregate market supply curve.


1
Expert's answer
2021-11-29T11:32:34-0500

a.      Determine the market aggregate supply function.

"Qseb=\u221264.5+37.5(10.68)\u22127.5(10)=261"



b..     Determine the inverse market supply function.

Holding all other things constant, the wage rate is constant at €10, so we have

"Qseb=\u221264.5+37.5Peb\u22127.5(10)=\u2212139.5+37.5Peb"

We now solve this for Peb:

"Peb = 3.72 + 0.0267Qeb"


c. Determine the slope of the aggregate market supply curve.


when Qeb rises by one unit, Peb rises by 0.0267 euros, so the slope of the supply curve is 0.0267, which is the coefficient on Qeb in the inverse supply function. Note that it is not 37.5.